# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ 7 A simulation study
We investigate the finite-sample performance of ML estimation for the model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$. We simulate 500 realizations of a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ for $d=2$, $\sigma^{2}=1$, $\kappa=0,1,2$, and $\mu=4,6,8$. The process is generated via Cholesky decomposition on a regular grid with spacing $0.025$ over $0,1]^{2}$ (1681 locations), followed by independent perturbations of each coordinate by a uniform random variable on [0.01,0.01][-0.01,0.01].
Throughout the study, $\kappa$ and $\mu$ are assumed known. We focus on ML estimation of $\sigma^{2}$, $a$, and the microergodic parameter $\sigma^{2}/a^{2\kappa+1}$ under both increasing- and fixed-domain scenarios. Following Zhang and Zimmerman (2005); Kaufman and Shaby (2013), these regimes can be mimicked by fixing the sampling grid and varying the spatial dependence. Accordingly, we consider $a=0.1,0.25,0.5$, interpreting $a=0.1$ as increasing-domain and $a=0.5$ as fixed-domain behavior.
Table 1 reports the bias and standard deviation (in parentheses) of the 500 ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$, for $i=1,\ldots,500$.
Table 1: Bias and standard deviation (in parenthesis) of ML estimates $\widehat{\sigma}_{i}^{2}$, $\widehat{a}_{i}$, and $\widehat{\sigma}_{i}^{2}/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ when $\kappa=0,1,2$, $\mu=4,6,8$, $a=0.1,0.25,0.5$ and $\sigma^{2}=1$.
@@
$\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$ $\mu=4$ $\mu=6$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ $0.0037$ $0.0025$ $0.0018$ $0.0060$ $0.0039$ $0.0028$ $0.0053$ $0.0038$ $0.0029$
$(0.0422)$ $(0.0382)$ $(0.0368)$ $(0.0485)$ $(0.0427)$ $(0.0398)$ $(0.0475)$ $(0.0432)$ $(0.0407)$
$a=0.25$ $0.0030$ $0.0047$ $0.0046$ $0.0085$ $0.0098$ $0.0095$ $0.0100$ $0.0116$ $0.0111$
$(0.0808)$ $(0.0622)$ $(0.0522)$ $(0.1082)$ $(0.0870)$ $(0.0737)$ $(0.1044)$ $(0.088)$ $(0.0773)$
$a=0.5$ 0.0023-0.0023 $0.0024$ $0.0045$ $0.0029$ $0.0079$ $0.0101$ $0.0052$ $0.0091$ $0.0116$
$(0.1446)$ $(0.1103)$ $(0.0900)$ $(0.1921)$ $(0.1592)$ $(0.135)$ $(0.1892)$ $(0.1599)$ $(0.1397)$
$\widehat{a}$ $a=0.1$ $0.0006$ $0.0007$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0002$ $0.0003$ $0.0003$
$(0.0049)$ $(0.0055)$ $(0.0068)$ $(0.0024)$ $(0.0028)$ $(0.0033)$ $(0.0017)$ $(0.0021)$ $(0.0026)$
$a=0.25$ $0.0007$ $0.0014$ $0.0016$ $0.0005$ $0.0008$ $0.0009$ $0.0004$ $0.0006$ $0.0007$
$(0.0214)$ $(0.0176)$ $(0.0157)$ $(0.0099)$ $(0.0086)$ $(0.0079)$ $(0.0059)$ $(0.0054)$ $(0.0052)$
$a=0.5$ 0.0015-0.0015 $0.0010$ $0.0023$ 0.0016-0.0016 0.0001-0.0001 $0.0008$ 0.0009-0.0009 0.0001-0.0001 $0.0005$
$(0.0737)$ $(0.0578)$ $(0.0488)$ $(0.0329)$ $(0.0280)$ $(0.0245)$ $(0.0196)$ $(0.0171)$ $(0.0154)$
@@
$\frac{\widehat{\sigma}^{2}}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.0150-0.0150 0.0219-0.0219 0.0151-0.0151 3.9684-3.9684 4.4120-4.4120 3.7659-3.7659 496.06-496.06 486.01-486.01 346.94-346.94
$(0.4370)$ $(0.5362)$ $(0.6792)$ $(51.3695)$ $(68.6746)$ $(88.9434)$ $(6397.8)$ $(8930.7)$ $(11628.7)$
$a=0.25$ $0.0045$ $0.0015$ 0.0017-0.0017 $0.0228$ 0.0470-0.0470 0.1218-0.1218 0.2553-0.2553 1.7130-1.7130 3.0083-3.0083
$(0.1499)$ $(0.1600)$ $(0.1708)$ $(2.5102)$ $(2.7402)$ $(2.9959)$ $(41.521)$ $(46.056)$ $(51.067)$
$a=0.5$ $0.0034$ $0.0033$ $0.0026$ $0.0113$ $0.0099$ $0.0065$ $0.0399$ $0.0319$ $0.0150$
$(0.0719)$ $(0.0740)$ $(0.0765)$ $(0.2923)$ $(0.3050)$ $(0.3187)$ $(1.1922)$ $(1.2484)$ $(1.3091)$
Overall, the ML estimators of $\sigma^{2}$ and $a$ are approximately unbiased across all configurations. As expected, their standard deviations increase when moving from the increasing-domain setting ($a=0.1$) to the fixed-domain setting ($a=0.5$). For fixed $\kappa$, larger values of $\mu$ generally reduce variability.
Regarding the microergodic parameter, under the increasing-domain scenario ($a=0.1$) the estimator exhibits moderate bias and relatively high variance, particularly for larger $\kappa$. Under the fixed-domain scenario ($a=0.5$), both bias and variance decrease substantially, and the sampling distribution appears closer to Gaussian, in agreement with the asymptotic theory. Moreover, the variability of the microergodic estimator increases with the smoothness parameter $\kappa$.
Figure 4 displays boxplots of the ML estimates of $\sigma^{2}$, $a$, and $\sigma^{2}/a^{2\kappa+1}$ for $\mu=4,6,8$ and $a=0.1$ (left) and $a=0.5$ (right), when $\kappa=0$. Under the increasing-domain scenario, the distributions of $\widehat{\sigma}^{2}$ and $\widehat{a}$ are approximately Gaussian. In the fixed-domain setting, the microergodic estimator becomes more concentrated around the true value and exhibits an approximately Gaussian distribution, as predicted by the theory.
@@
($a=0.1$) ($a=0.5$)
Figure 4: Left part: Boxplots of ML estimates of $\sigma^{2}$, $a$ and the microergodic parameter $\frac{\sigma^{2}}{a^{2\kappa+1}}$ (from left to right) when estimating the covariance model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$ with $\sigma^{2}=1$, for $\mu=4,6,8$ and $a=0.1$ (increasing domain scenario). Right part: the same Figure with $a=0.5$ (fixed domain scenario). In both cases $\kappa=0$.
@@ 8 A real data application
For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$, and the proposed ${\cal H}_{\kappa,\mu,\beta,2}$. Here $\kappa$ denotes smoothness, $\beta$ the scale parameter, and $\mu$ the shape parameter.
@@
Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values. Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model. Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$ achieve the best AIC values.
Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,\beta,2}$ correlation. The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,\beta}$, ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$. Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
@@
${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$ ${\cal MT}_{\kappa,\beta}$ ${\cal GW}_{\kappa,\mu,\beta}$ ${\cal H}_{\kappa,\mu,\beta,2}$
@@
We assess predictive performance via repeated cross-validation. For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction. Table 2 reports the average RMSE and MAE. The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,\beta}$ and ${\cal H}_{\kappa,\mu,\beta,2}$.
@@ 9 Concluding Remarks
In this paper we study the Gauss hypergeometric covariance model, a flexible class introduced by Emery and Alegría (2022) that encompasses both the Matérn and Generalized Wendland families. We derive necessary and sufficient validity conditions and show that the model exhibits structural identifiability issues under both increasing- and fixed-domain asymptotics. To address this issue, we introduce a parsimonious compactly supported subclass selected through an optimality criterion based on maximum integral range. In this sense, the proposed hypergeometric model can be viewed as a structural refinement of the Generalized Wendland construction.
Although simple closed-form expressions are not generally available for nonnegative integer smoothness parameters, efficient numerical evaluation of the Gauss hypergeometric function is widely supported in modern scientific computing libraries. In particular, reliable implementations are available in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python. The GeoModels package (Bevilacqua et al., 2026), used in this paper for numerical experiments and applications, computes the proposed model via the SciPy implementation of the hypergeometric function (Virtanen et al., 2020).

### v3

@@ 7 A simulation study
We study the finite-sample behavior of ML estimation under the proposed model $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,d}$. We generate $500$ independent realizations of a zero-mean Gaussian random field on $0,1]^{2}$ ($d=2$) with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, and $\mu\in\{4,8\}$. Each realization is simulated via Cholesky decomposition at $n=2500$ locations randomly sampled uniformly over the unit square.
Throughout, we treat $\kappa$ as known and focus on ML estimation of $\sigma^{2}$, $a$, $\mu$, and the microergodic quantity $\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$. To create a gradual transition from increasing-domain to fixed-domain behavior while keeping the sampling design fixed, we follow Zhang and Zimmerman (2005); Kaufman and Shaby (2013) and vary the scale parameter $a$. Specifically, we consider $a\in\{0.1,0.3,0.6,0.9\}$. Here $a=0.1$ corresponds to weak dependence relative to the unit-square domain and thus approximates an increasing-domain regime, whereas $a=0.9$ induces strong dependence and is much closer to a fixed-domain setting.
Table 1 reports the empirical bias and standard deviation (in parentheses) of the $500$ ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$, $\widehat{\mu}_{i}$ and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $i=1,\ldots,500$. The main pattern is clear: the individual covariance parameters $\sigma^{2}$, $a$, and $\mu$ become more variable as $a$ increases, whereas the microergodic parameter behaves in the opposite way.
Table 1: Bias and standard deviation (in parentheses) of ML estimates $\widehat{\sigma}^{2}_{i}$, $\widehat{a}_{i}$$\widehat{\mu}_{i}$, and $\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$ for $i=1,\ldots,500$ when estimating a zero-mean Gaussian RF with covariance $\sigma^{2}\mathcal{H}_{\kappa,\mu,a,2}$ on $0,1]^{2}$ ($n=2500$ locations), with $\sigma^{2}=1$, $\kappa\in\{0,1,2\}$, $\mu\in\{4,8\}$, and $a\in\{0.1,0.3,0.6,0.9\}$.
@@
$\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$ $\mu=4$ $\mu=8$
$\widehat{\sigma}^{2}$ $a=0.1$ 0.00299 0.00107 0.00433 0.00184 0.00408 0.00212
(0.04138) (0.03205) (0.04914) (0.03759) (0.04794) (0.03889)
$a=0.3$ 0.01575 0.00504 0.01504 0.00770 0.01173 0.00726
(0.10407) (0.05976) (0.13271) (0.08844) (0.12980) (0.09514)
$a=0.6$ 0.06435 0.01181 0.04389 0.01385 0.03073 0.01553
(0.26849) (0.11312) (0.25614) (0.16863) (0.23655) (0.18127)
$a=0.9$ 0.12097 0.02437 0.09481 0.01848 0.05189 0.02400
(0.38703) (0.17580) (0.41788) (0.24205) (0.32244) (0.25633)
$\widehat{a}$ $a=0.1$ 0.02090 0.08978 0.00276 0.01635 0.00137 0.00612
(0.10465) (0.20471) (0.01397) (0.06865) (0.00827) (0.02363)
$a=0.3$ 0.81440 0.32342 0.02153 0.05872 0.00511 0.01092
(2.75327) (0.75816) (0.10263) (0.26379) (0.04194) (0.06661)
$a=0.6$ 3.91933 2.29317 0.55575 0.45367 0.04601 0.07026
(8.37251) (4.53301) (2.64211) (1.50576) (0.22104) (0.33102)
$a=0.9$ 7.46177 4.50776 1.66114 1.55866 0.16050 0.22433
(14.27313) (7.73620) (5.73980) (4.14495) (0.77112) (0.80880)
$\widehat{\mu}$ $a=0.1$ 1.00129 8.12410 0.17091 1.81217 0.10804 0.81383
(5.12648) (18.51370) (0.88371) (7.52792) (0.66348) (3.13054)
$a=0.3$ 12.17228 9.21399 0.42205 1.98131 0.12180 0.42181
(41.02159) (21.40586) (2.01811) (8.89166) (0.96416) (2.54729)
$a=0.6$ 29.22591 32.41524 5.42300 7.57716 0.55245 1.34734
(61.35913) (63.46363) (25.61297) (24.98706) (2.54928) (6.27535)
$a=0.9$ 36.69791 42.83158 10.93102 17.38350 1.29228 2.84537
(68.21306) (72.15506) (37.54848) (45.95964) (6.09577) (10.03183)
$\frac{\widehat{\sigma}^{2}F(\kappa,\mu,d)}{\widehat{a}^{2\kappa+1}}$ $a=0.1$ 0.02847 -0.15330 567.68210 7656.58100 $1.40740\times 10^{7}$ $4.32344\times 10^{8}$
(1.79328) (5.08178) (12695.11000) (107860.90000) ($2.18681\times 10^{8}$) ($3.79118\times 10^{9}$)
$a=0.3$ 0.00880 -0.04215 -0.59747 -0.93738 -2994.48500 -48566.15000
(0.45226) (0.95325) (296.09960) (1651.58800) (448665.40000) (4813284.00000)
$a=0.6$ 0.00331 -0.01499 0.66440 -0.59339 165.89200 -325.76500
(0.21584) (0.43426) (32.70379) (167.22660) (11788.42000) (115466.10000)
$a=0.9$ 0.00148 -0.01088 0.05882 -0.61039 -4.95065 -62.59252
(0.14205) (0.28237) (9.47927) (46.79391) (1482.45500) (13819.55000)
This is especially evident for $\widehat{a}$ and $\widehat{\mu}$. For instance, when $\kappa=0$ and $\mu=4$, the empirical standard deviation of $\widehat{a}$ increases from $0.10465$ at $a=0.1$ to $2.75327$ at $a=0.3$, then to $8.37251$ at $a=0.6$, and to $14.27313$ at $a=0.9$; the standard deviation of $\widehat{\mu}$ rises from $5.12648$ to $41.02159$, then to $61.35913$, and finally to $68.21306$. The estimator $\widehat{\sigma}^{2}$ is more stable, but it follows the same qualitative trend.
These results are fully consistent with fixed-domain asymptotics: the individual covariance parameters are not consistently estimable, and stronger dependence effectively reduces the amount of independent information. The effect of $\mu$ is visible but secondary relative to the role of $a$. In several configurations, $\mu=8$ leads to somewhat larger variability than $\mu=4$, especially for $\widehat{\mu}$ and in the stronger-dependence regimes. Likewise, the overall inferential difficulty increases with $\kappa$, particularly for quantities involving the scale parameter.
The microergodic estimator behaves in the opposite way. For
$\widehat{\theta}=\frac{\widehat{\sigma}^{2}F(\kappa,\widehat{\mu},d)}{\widehat{a}^{2\kappa+1}},$
both bias and dispersion decrease markedly as $a$ increases. This is most striking for larger $\kappa$. For instance, when $\kappa=2$ and $\mu=4$, the empirical standard deviation of $\widehat{\theta}$ drops from $2.18681\times 10^{8}$ at $a=0.1$, to $4.48665\times 10^{5}$ at $a=0.3$, to $1.17884\times 10^{4}$ at $a=0.6$, and to $1482.45500$ at $a=0.9$. Even for $\kappa=0$, where $\widehat{\theta}$ is already much more stable, the same monotone improvement is evident.
This contrast is the central empirical message of the simulation study. As the regime moves from increasing-domain-like to fixed-domain-like behavior, the individual covariance parameters deteriorate, while the microergodic parameter improves. The distinction becomes sharper as $\kappa$ increases, since the factor $a^{-(2\kappa+1)}$ amplifies scale estimation error in weak-dependence settings, but defines the identifiable combination under fixed-domain asymptotics.
Figure 4 shows normal QQ-plots for the standardized microergodic estimates in the strong-dependence setting ($a=0.9$) with $\kappa=1$, comparing $\mu=4$ (left) and $\mu=8$ (right). Specifically, we plot
@@
$Z_{i}\;=\;\frac{\sqrt{n}\,\big(\widehat{\theta}_{i}-\theta_{0}\big)}{\sqrt{2}\,\theta_{0}},\qquad\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1},\quad\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1},$
for $i=1,\dots,500$. In both cases, the QQ-plots align reasonably well with the standard normal reference line, supporting the relevance of the fixed-domain Gaussian approximation at this sample size.
Overall, the simulations strongly support the theoretical results of Section 6: the individual covariance parameters are increasingly unstable as dependence strengthens, whereas the microergodic parameter shows the opposite trend and is the meaningful inferential target under fixed-domain asymptotics.
@@
($\mu=4$) ($\mu=8$)
Figure 4: Normal $Q$-plots of standardized microergodic estimates in the fixed-domain setting ($a=0.9$) with $\kappa=1$ and $\sigma^{2}=1$, comparing $\mu=4$ (left) and $\mu=8$ (right). The standardized statistics are $Z_{i}=\sqrt{n}\,(\widehat{\theta}_{i}-\theta_{0})/(\sqrt{2}\,\theta_{0})$, where $\widehat{\theta}_{i}=\widehat{\sigma}^{2}_{i}F(\kappa,\widehat{\mu}_{i},d)/\widehat{a}_{i}^{2\kappa+1}$, $\theta_{0}=\sigma^{2}F(\kappa,\mu,d)/a^{2\kappa+1}$, and $n=2500$.
@@ 8 A real data application
For the correlation of $Z(\mathbf{s})$, we compare three models: ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$, and the proposed ${\cal H}_{\kappa,\mu,a,2}$. Here $\kappa$ denotes smoothness, $a$ the scale parameter, and $\mu$ the shape parameter.
@@
Table 2 reports ML estimates, associated standard errors, maximized log-likelihood values, and AIC values. Overall, the AIC criterion favors the Tukey-$h$ specification, irrespective of the correlation model. Among the correlation structures, the compactly supported ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$ achieve the best AIC values.
Figure 6 compares the normalized histogram of the residuals with the fitted density (left panels) and the empirical semivariogram with the fitted semivariogram (right panels), for the Gaussian (first row) and Tukey-$h$ (second row) models using the estimated ${\cal H}_{\kappa,\mu,a,2}$ correlation. The Tukey-$h$ RF provides a noticeably better fit to both marginal tail behavior and spatial dependence.
Table 2: ML estimates with associated standard error (in parentheses), RMSE, MAE and percentage of zero values in the covariance matrix for the Gaussian and Tukey-$h$ RFs with underlying correlation functions ${\cal MT}_{\kappa,a}$, ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$. Last row reports the time (in seconds) needed to compute the Cholesky factor of the covariance matrix.
@@
${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$ ${\cal MT}_{\kappa,a}$ ${\cal GW}_{\kappa,\mu,a}$ ${\cal H}_{\kappa,\mu,a,2}$
@@
We assess predictive performance via repeated cross-validation. For each of 100 repetitions, 85% of locations are used for parameter estimation and the remaining 15% for prediction. Table 2 reports the average RMSE and MAE. The best predictive performance is achieved by the Tukey-$h$ RF combined with compactly supported correlations ${\cal GW}_{\kappa,\mu,a}$ and ${\cal H}_{\kappa,\mu,a,2}$.
@@ 9 Concluding Remarks
We revisited the Gauss hypergeometric covariance class (Emery and Alegría, 2022), a flexible family that includes the celebrated Matérn and generalized Wendland models as special cases. We first refined its validity conditions, deriving sharp constraints over a substantial region of the parameter space and pinpointing where a gap between necessary and sufficient conditions still remains.
We then clarified that the class features two distinct non-identifiability mechanisms: a structural one, stemming from an intrinsic symmetry in the shape parameters, and the classical fixed-domain one, whereby different parameter values can induce equivalent Gaussian measures and only microergodic combinations are consistently estimable. The structural ambiguity is removed by an ordering constraint, yielding a unique representation on the restricted space.
Within this identifiable formulation, we introduced a parsimonious compactly supported subclass, $\mathcal{H}$, selected via a maximum integral range criterion. This choice provides a canonical representative with maximal low-frequency power among admissible models with fixed smoothness and support. The resulting family generalizes the classical spherical model (Chilès and Delfiner, 2012) while retaining continuous smoothness control, compact support, and simple necessary and sufficient validity conditions. We also derived scale-mixture representations and explicit links to the Matérn model through compact-support reparameterizations.
From an inferential standpoint, we developed fixed-domain asymptotics for likelihood-based estimation under $\mathcal{H}$. For fixed smoothness, we proved consistency and asymptotic normality for the microergodic-parameter estimator, and established stability when nuisance parameters are estimated by ML over compact sets. Simulations and a climate-data application indicate that $\mathcal{H}$ can offer a competitive compromise between Matérn and generalized Wendland models, combining accuracy with sparse covariance matrices and notable computational gains, for both Gaussian and Tukey-hh random fields.
Several directions remain open, including closing the remaining validity gap in Theorem 2, extending fixed-domain theory to unknown smoothness, and developing scalable alternatives to full ML. Finally, anisotropic, nonstationary, and multivariate extensions would further broaden the practical scope of the hypergeometric framework.
Although closed forms are available only in special cases, efficient numerical evaluation of Gauss hypergeometric functions is widely supported, including in the GNU Scientific Library (Gough, 2009) and in major environments such as R, MATLAB, and Python. The GeoModels package (Bevilacqua et al., 2026) used in this paper computes the proposed model via the SciPy implementation (Virtanen et al., 2020).